What is the kinetic energy of automobile with a mass of 1.25 x 10^3 kg traveling at a speed of 10.0m/s


1
Expert's answer
2021-04-15T07:17:30-0400

The kinetic energy of the automobile is:

"E = \frac{m v^2}{2} = \frac{1.25 \cdot 10^3 kg \cdot 10^2 \frac{m^2}{s^2}}{2} = 62500 J = 62.5 kJ."